OpenStudy (anonymous):

Use the Laws of Exponents to simplify the 4 OVER 40 and 4 over 20 expression Explain how you found your answer and write the answer as an exponent

OpenStudy (anonymous):

so? \[4=2^2=2*2\] factor out 40 and 20 40 = 2*2*2*5 20=2*2*5 and simplify them?

OpenStudy (anonymous):

40=40 20=20

OpenStudy (anonymous):

2 x 2 x 2 x 5 =? and 2 x 2 x 5 =? \[2^2 * 2 *5 =?\] \[2^2 * 5=?\]

OpenStudy (anonymous):

20

OpenStudy (anonymous):

we know that the number \[2^3=8\] and you mutiply it by 5 to make 40 then we have\[4=2^2\] and we devide \[2^2 / (2^3*5)\] then you'l have 1/10

OpenStudy (anonymous):

the second you follow the same method